Problem 1.

Resistor for electronic circuits are manufactured on a high-speed automated machine. The machine is set up to produce a large run of resistors of 1,000 ohms each.

To set up the machine and to create a control chart to be used throughout the run, 15 samples were taken with four resistors in each sample. The complete list of samples and their measured values are as follows:

Sample #

Readings (in ohms)

1

1010

991

985

986

2

995

996

1009

994

3

990

1003

1015

1008

4

1015

1020

1009

998

5

1013

1019

1005

993

6

994

1001

994

1005

7

989

992

982

1020

8

1001

986

996

996

9

1006

989

1005

1007

10

992

1007

1006

979

11

996

1006

997

989

12

1019

996

991

1011

13

981

991

989

1003

14

999

993

988

984

15

1013

1002

1005

992

  1. (10 points) Calculate control limits for R-chart.
  2. (10 points) Calculate control limits for x̄ - chart.
  3. (20 points) From the charts, what comments can you make about the process? and why?
Answer #1

Formula

a)

R-chart

UCL = D4*R-bar [For sample size 4, D4 = 2.282]
=2.282*21.73333333 = 49.59546666


LCL = D3*R-bar [For sample size 4, D3 = 0] = 0

b)

x-bar-bar chart

UCL = x-bar-bar+A2.R-bar [For sample size 4, A2 = 0.729]

=999.1+0.729*21.73333333

=1014.9436

LCL = x-bar-bar-A2.R-bar [For sample size 4, A2 = 0.729]

=999.1-0.729*21.73333333

= 983.2564

c)

The process is in control in terms of R-chart and x-bar-bar chart as all the sample means and ranges are within control limits